OpenStudy (anonymous):

A pet store owner found that 1/8 of his 640 pets are cats and 1/10 of the remaining animals are dogs, How many of the pets are dogs?

OpenStudy (anonymous):

640/8 = 80 cats 640-80=560 560/10 = 56 56 dogs

OpenStudy (anonymous):

The remaining of animals would be 560 since you already know that there 80 cats. So you just divide by 10 from the remaining to get the amount of dogs. which therefore is 56